### v4

# Making AI Less “Thirsty”: Uncovering and Addressing the Secret Water Footprint of AI Models

Abstract
The growing carbon footprint of artificial intelligence (AI) has been undergoing public scrutiny. Nonetheless, the equally important water (withdrawal and consumption) footprint of AI has largely remained under the radar. For example, training the GPT-3 language model in Microsoft’s state-of-the-art U.S. data centers can directly evaporate 700,000 liters of clean freshwater, but such information has been kept a secret. More critically, the global AI demand is projected to account for 4.2 – 6.6 billion cubic meters of water withdrawal in 2027, which is more than the total annual water withdrawal of 4 – 6 Denmark or half of the United Kingdom. This is concerning, as freshwater scarcity has become one of the most pressing challenges. To respond to the global water challenges, AI can, and also must, take social responsibility and lead by example by addressing its own water footprint. In this paper, we provide a principled methodology to estimate the water footprint of AI, and also discuss the unique spatial-temporal diversities of AI’s runtime water efficiency. Finally, we highlight the necessity of holistically addressing water footprint along with carbon footprint to enable truly sustainable AI.
1 Introduction
$\bullet$ “Water is a finite resource, and every drop matters.” — Facebook (now Meta) Sustainability Report 2020 [1]. $\bullet$ “Fresh, clean water is one of the most precious resources on Earth … Now we’re taking urgent action to support water security and healthy ecosystems.” — Google’s Water Commitment 2023 [2]. $\bullet$ “Water is a human right and the common development denominator to shape a better future. But water is in deep trouble.” — U.N. Secretary-General António Guterres at the U.N. Water Conference 2023 [3]. $\bullet$ “Historic droughts threaten our supply of water … As the source of both life and livelihoods, water security is central to human and national security.” — The U.S. White House Action Plan on Global Water Security 2022 [4].
Artificial intelligence (AI) has enabled remarkable breakthroughs in numerous areas of critical importance, including tackling global challenges such as climate change. On the other hand, many AI models, especially large generative ones like GPT-4, are trained and deployed on energy-hungry servers in warehouse-scale data centers, accelerating the data center energy consumption at an unprecedented rate [5]. As a result, AI’s carbon footprint has been undergoing scrutiny, driving the recent progress in AI carbon efficiency [6, 7]. However, AI’s water footprint — many millions of liters of freshwater consumed for cooling the servers and for electricity generation — has largely remained under the radar and keeps escalating. If not properly addressed, AI’s water footprint can potentially become a major roadblock to sustainability and create social conflicts as freshwater resources suitable for human use are extremely limited and unevenly distributed.
As acknowledged in Google’s sustainability report [8] and the recent U.S. Department of Energy study [5], the expansion of AI products and services is a key driver of the rapid increase in data center water consumption. Even excluding the water usage in leased third-party colocation facilities, Google’s self-owned data centers alone directly withdrew 29 billion liters and consumed (i.e., evaporated) more than 23 billion liters of freshwater for on-site cooling in 2023, nearly 80% of which was potable water [8]. [^2] The detailed difference between water withdrawal and water consumption is presented in Section 2.1. This amount of annual water consumption even rivals that of a major household-name beverage company [9]. Importantly, Google’s data center water consumption increased by $\sim$20from 2021 to 2022 and by $\sim$17from 2022 to 2023 [8], and Microsoft saw $\sim$34and $\sim$22increases over the same periods, respectively [10]. Furthermore, according to the U.S. Department of Energy, the total annual on-site water consumption by U.S. data centers in 2028 could double or even quadruple the 2023 level, reaching approximately 150 – 280 billion liters and further stressing the water infrastructures [5].
AI represents the fastest expanding workloads in data centers [8, 5]. For example, a recent study suggests that the global AI could consume 85 – 134 TWh of electricity in 2027 [11], whereas a more aggressive projection by the U.S. Department of Energy predicts that AI servers’ electricity consumption in the U.S. alone will surpass 150 – 300 TWh in 2028 [5]. Even considering the lower estimate, the combined scope-1 and scope-2 water withdrawal of global AI is projected to reach 4.2 – 6.6 billion cubic meters in 2027, which is more than the total annual water withdrawal of 4 – 6 Denmark or half of the United Kingdom. [^3] The scope definition of water usage [12] is in line with that of carbon emissions and is discussed in Section 2.2. Our scope-2 water withdrawal (and consumption when applicable) is for location-based electricity generation throughout the paper. Large data centers often adopt sustainability programs (e.g., renewable purchasing agreements) to offset their location-based electricity usage and thus may have lower market-based carbon and water footprints. Simultaneously, a total of 0.38 – 0.60 billion cubic meters of water will be evaporated and considered “consumption” due to the global AI demand in 2027. Moreover, these global estimates will be exceeded by the total water withdrawal and consumption attributed to AI in the U.S. alone in 2028 if the U.S. Department of Energy’s projection comes to fruition.
Despite its profound environmental and societal impact, the increasing water footprint of AI has received disproportionately less attention from the AI community as well as the general public. For example, while the scope-2 carbon emissions are routinely included as part of AI model cards, even scope-1 direct water usage (either withdrawal or consumption) is missing, let alone scope-2 water usage. This may impede innovations to enable water sustainability and build truly sustainable AI. Crucially, water and carbon footprints are complementary to, not substitutable of, each other for understanding the environmental impacts. Indeed, optimizing for carbon efficiency does not necessarily result in, and may even worsen, water efficiency, which varies with the fuel mixes for electricity generation and outside weather in a unique way [13, 5].
To ensure that the growth in AI does not exacerbate the global water stresses or outweigh the environmental benefits it provides, it is critical to uncover and address AI’s hidden water footprint amid the increasingly severe freshwater scarcity crisis, worsened extended droughts, and quickly aging public water infrastructure. The urgency can also be reflected in part by the recent commitment to “Water Positive by 2030” from industry leaders, including Google [8] and Microsoft [10], by policy guidelines and legislative efforts to mitigate AI’s water consumption [14, 15], and by the inclusion of water consumption as a key environmental metric into the first international standard on sustainable AI to be published by the ISO/IEC [16]. Moreover, setting targets for minimizing water consumption is included as part of the recent U.S. executive order for advancing AI data center infrastructures [17].
In this paper, we advocate for a holistic approach to sustainable AI that extends beyond the carbon footprint to also address the water footprint. Specifically, we present a principled methodology to estimate AI’s total water footprint, including both operational water and embodied water. By taking the GPT-3 model with 175 billion parameters as an example [18], we show that training GPT-3 in Microsoft’s U.S. data centers can consume a total of 5.4 million liters of water, including 700,000 liters of scope-1 on-site water consumption. Additionally, GPT-3 needs to “drink” (i.e., consume) a 500ml bottle of water for roughly 10 – 50 medium-length responses, depending on when and where it is deployed. Note that our estimate of inference water consumption for GPT-3 is conservative, and the actual water consumption could be several times higher.
Next, we emphasize the need for increasing transparency of AI’ water footprint, including disclosing more information about operational data and keeping users informed of the runtime water efficiency. We show that WUE (Water Usage Effectiveness, a measure of water efficiency) varies both spatially and temporally, suggesting that “when” and “where” to perform AI training can significantly affect the water footprint. Finally, we highlight the necessity of holistically addressing the water footprint along with the carbon footprint to enable truly sustainable AI — the water footprint of AI can no longer stay under the radar.
2 Background
2.1 Water Withdrawal vs. Water Consumption
There are two related but different concepts — water withdrawal and water consumption, both of which are important for understanding the impacts on water stress and availability [19, 20].
$\bullet$ Water withdrawal: It refers to freshwater taken from the ground or surface water sources, either temporarily or permanently, and then used for agricultural, industrial, or municipal uses (normally excluding water used for hydroelectricity generation) [19]. As water is a finite shared resource, water withdrawal indicates the level of competition as well as dependence on water resources among different sectors.
$\bullet$ Water consumption: It is defined as “water withdrawal minus water discharge”, and means the amount of water “evaporated, transpired, incorporated into products or crops, or otherwise removed from the immediate water environment” [20]. Water consumption reflects the impact on downstream water availability and is crucial for assessing watershed-level scarcity [19].
These two types of water usage correspond to two different water footprints, i.e., water withdrawal footprint (WWF) [21, 12] and water consumption footprint (WCF), respectively [22]. By default, water footprint refers to the water consumption footprint unless otherwise specified.
2.2 How Does AI Use Water?
AI’s water usage spans three scopes: on-site water for data center cooling (scope 1), off-site water for electricity generation (scope 2), and supply-chain water for server manufacturing (scope 3).
2.2.1 Scope-1 Water Usage
Nearly all the server energy is converted into heat, which must then be removed from the data center server room to avoid overheating. This process involves two sequential stages: server-level cooling followed by facility-level cooling.
In the server-level cooling stage, heat is transferred from the servers to the facility or a heat exchanger, typically using either air or liquid cooling methods (e.g., direct-to-chip cooling or immersion cooling), which do not evaporate or consume water. In general, new data centers dedicated to AI training often rely on liquid cooling due to the high server power densities.
In the facility-level cooling stage, heat is rejected from the data center facility to the outside environment. While there are various cooling methods, water-intensive cooling towers and water evaporation-assisted air cooling are two common approaches used in many data centers, including those operated by major technology companies [8, 5].
Figure 1: An example of data center’s operational water usage: on-site scope-1 water usage for data center cooling (via cooling towers in the example), and off-site scope-2 water usage for electricity generation. The icons for AI models are only for illustration purposes.
Cooling tower. As illustrated in Figure 1, some water is evaporated (i.e., “consumed”) in the cooling tower to dissipate heat into the environment, while the remaining water moves along an open loop to the heat exchanger to further absorb the server heat. Additionally, non-evaporated water can be recycled only a few times (typically 3–10 cycles, depending on water quality) before discharge, requiring continuous clean freshwater replenishment to prevent mineral and salt buildup. Thus, to keep the cooling tower working, new water must be constantly added to make up for the evaporated water and discharged water. Importantly, clean freshwater (potable water in many cases [8]) is needed to avoid pipe clogs and/or bacterial growth.
For cooling towers, water withdrawal refers to the amount of added water, including both evaporated water and discharged water, while water consumption exclusively indicates the amount of evaporated water. With good water quality, roughly 80% of water withdrawal is evaporated and considered “consumption” [8]. On average, depending on the weather conditions and operational settings, data centers can evaporate approximately 1 – 9 liters per kWh of server energy: 1 L/kWh for Google’s annualized global on-site water efficiency [8] and 9 L/kWh for a large commercial data center during the summer in Arizona [23].
Air cooling with water evaporation assistance. When the climate condition is appropriate, data centers may use “free” outside air to directly reject the heat to the outside environment. Nonetheless, water evaporation is still needed when the outside air is too hot (e.g., higher than 85 degrees Fahrenheit); additionally, water is also needed for humidity control when the outside air is too dry [24]. The added water is considered “withdrawal”, out of which about 70% is consumed based on Meta’s report [25]. Generally, outside air cooling is more water-efficient than cooling towers on average. However, hot weather raises the evaporative water demand and maximum water consumption, potentially stressing local water supplies during peak demand on hot days. Additionally, the application of outside air cooling may have challenges in hot regions and/or for many colocation facilities that are located in business districts.
Some data centers may opt for dry coolers, which consume no on-site water year-round [26]. However, this approach typically increases cooling energy consumption compared to water-based cooling methods, potentially exacerbating the overall stress on water resources due to higher scope-2 water consumption.
2.2.2 Scope-2 Water Usage
In many countries, thermoelectric power is among the top sectors in terms of water withdrawal and water consumption [12]. Thus, similarly to scope-2 carbon emissions, data centers are accountable for off-site scope-2 water usage associated with electricity consumption, which forms part of the “true water cost of data centers,” as highlighted by the U.S. Department of Energy [5].
Different power plants use different amounts of water for each kWh generation, depending on the cooling techniques. Typically, water withdrawal due to hydropower generation is excluded, but water consumption due to increased water evaporation rates from hydropower generation is included [5]. For electricity generation, the U.S. national average water withdrawal and consumption are estimated at about 43.8 L/kWh [27] and 3.1 L/kWh [12], respectively. Meta’s self-reported scope-2 water consumption for its global data center fleet was 3.7 L/kWh (i.e., 55,475 megaliters divided by 14,975,435 MWh) in 2023 [25].
2.2.3 Scope-3 Water Usage
AI chip and server manufacturing uses a huge amount of water [28, 29]. For example, ultrapure water is needed for wafer fabrication and water is also needed for keeping semiconductor plants cool. Importantly, the discharged water may contain toxic chemicals and/or hazardous wastes. While water recycling at semiconductor plants can effectively reduce water withdrawal, the recycling rate in many cases remains low, e.g., the average recycling rate for wafer plants and semiconductor plants in Singapore are 45% and 23%, respectively [29]. Although largely obscure, scope-3 water usage is likely significant [28]. For instance, Apple reports that its supply chain accounts for 99% of its total water footprint [30].
It is important to recognize that, unlike agriculture whose water footprint is mostly green (i.e., water stored in soil and used by plants), the majority of AI’s water footprint is blue water extracted from rivers, lakes, or groundwater, which is directly accessible for human use but often more limited in availability.
3 Estimating AI’s Water Footprint
We present a general methodology for estimating AI’s water consumption footprint. To obtain the water withdrawal footprint, we simply replace the WUE with water withdrawal efficiency.
3.1 Operational Water Footprint
We collectively refer to on-site scope-1 water and off-site scope-2 water as the operational water.
$\bullet$ On-site WUE. We denote the on-site scope-1 WUE at time $t$ by $\rho_{{s1},t}$, which is defined as the ratio of the on-site water consumption to server energy consumption and varies over time depending on the outside temperature (see [13] for an example of on-site WUE based on cooling towers). Concretely, $\rho_{{s1},t}$ increases significantly for cooling towers when the outside wet bulb temperature increases, and increases for outside air cooling when the outside dry bulb temperature is too hot or the humidity is too low.
$\bullet$ Off-site WUE. We denote the off-site scope-2 WUE at time $t$ as $\rho_{s2,t}$, which is defined as the ratio of off-site water consumption for each kWh of electricity consumption and measures the electricity water intensity factor (EWIF). While there are different methods to estimate $\rho_{s2,t}$, a common one is weighted averaging: $\rho_{s2,t}=\frac{\sum_{k}{b_{k,t}}\times EWIF_{k}}{\sum_{k}{b_{k,t}}}$ where $b_{k,t}$ denotes the amount of electricity generated from fuel type $k$ at time $t$ for the grid serving the data center under consideration, and $EWIF_{k}$ is the EWIF for fuel type $k$ [31, 32]. Thus, variations in energy fuel mixes of electricity generation result in temporal variations of the off-site WUE. Moreover, the off-site WUE also varies across regions due to different energy fuel mixes [12, 5].
$\bullet$ Operational water footprint. Consider a time-slotted model $t=1,2,\cdots,T$, where the length of each time slot depends on how frequently we want to assess the operational water footprint. At time $t$, suppose that an AI model uses energy $e_{t}$ which can be measured using power meters and/or servers’ built-in tools, and the data center hosting the AI model has a power usage effectiveness (PUE) of $\theta_{t}$ that accounts for the non-IT energy overhead. Then, the total operational water footprint of the AI model can be written as $WaterOperational=\sum_{t=1}^{T}e_{t}\cdot\left[\rho_{s1,t}+\theta_{t}\cdot\rho_{s2,t}\right]$.
3.2 Embodied Water Footprint
Similar to accounting for the embodied carbon footprint [33], the total scope-3 water footprint is amortized over the lifespan of a server. Specifically, if $W$ represents the total water used to manufacture the AI servers and the servers are expected to operate for a period of $T_{0}$, then the embodied water footprint over a period of $T$ is calculated as $WaterEmbodied=\frac{T\cdot W}{T_{0}}$
By adding up the operational and embodied water footprints, we obtain the total water footprint as $WaterTotal=\sum_{t=1}^{T}e_{t}\cdot\left[\rho_{s1,t}+\theta_{t}\cdot\rho_{s2,t}\right]+\frac{T\cdot W}{T_{0}}$. In practice, to obtain a rough estimate, we can use the average values for the annualized WUE and the estimated AI server energy consumption.
3.3 Case Study: Estimating GPT-3’s Operational Water Consumption Footprint
Table 1: Estimate of GPT-3’s operational water consumption footprint. “*” denotes data centers under construction as of July 2023, whose PUE and WUE are projected by Microsoft.
Location PUE On-site WUE (L/kWh) Off-site EWIF (L/kWh) Water for Training (million L) Water for Each Request (mL) # of Requests  for 500ml  Water
On-site
Water
Off-site
Water
Total
Water
On-site
Water
Off-site
Water
Total
Water
U.S. Average 1.170 0.550 3.142 0.708 4.731 5.439 2.200 14.704 16.904 29.6
Arizona 1.180 1.630 4.959 2.098 7.531 9.629 6.520 23.406 29.926 16.7
Georgia* 1.120 0.060 2.309 0.077 3.328 3.406 0.240 10.345 10.585 47.2
Illinois 1.350 0.740 2.233 0.952 3.880 4.833 2.960 12.060 15.020 33.3
Iowa 1.160 0.140 3.104 0.180 4.634 4.814 0.560 14.403 14.963 33.4
Texas 1.280 0.250 1.287 0.322 2.120 2.442 1.000 6.590 7.590 65.9
Virginia 1.140 0.140 2.385 0.180 3.499 3.679 0.560 10.875 11.435 43.7
Washington 1.150 0.950 9.501 1.223 14.063 15.285 3.800 43.706 47.506 10.5
Wyoming 1.110 0.130 2.574 0.167 3.677 3.845 0.520 11.429 11.949 41.8
Australia* 1.120 0.012 4.259 0.015 6.138 6.154 0.048 19.078 19.126 26.1
Denmark* 1.160 0.010 3.180 0.013 4.747 4.760 0.040 14.754 14.794 33.8
Finland* 1.120 0.010 4.542 0.013 6.548 6.561 0.040 20.350 20.390 24.5
India* 1.430 0.000 3.445 0.000 6.340 6.340 0.000 19.704 19.704 25.4
Indonesia* 1.320 1.900 2.271 2.445 3.858 6.304 7.600 11.992 19.592 25.5
Ireland 1.190 0.020 1.476 0.026 2.261 2.287 0.080 7.027 7.107 70.4
Mexico* 1.120 0.056 5.300 0.072 7.639 7.711 0.224 23.742 23.966 20.9
Netherlands 1.140 0.060 3.445 0.077 5.054 5.131 0.240 15.708 15.948 31.4
Sweden 1.160 0.090 6.019 0.116 8.986 9.101 0.360 27.927 28.287 17.7
The core of ChatGPT, a popular online service, is a large language model (LLM) based on subsequent versions of GPT-3. We present a case study to estimate the operational water consumption for the full GPT-3 model with 175 billion parameters [18]. We exclude embodied water footprint due to the lack of public data for scope-3 water usage. We choose GPT-3 as Microsoft publishes its location-wise WUE and PUE [34, 35]. The results are summarized in Table 1.
3.3.1 Training
GPT-3 was trained and deployed by OpenAI in Microsoft’s data centers, with an estimated training energy of 1287 MWh [36]. In line with the practice of estimating the carbon footprint, we use the most recent annualized average on-site PUE and WUE for each location, as reported by Microsoft [34, 35]. For power plant water efficiency, different references may provide different estimates of EWIF. Thus, for consistency across regions, we use the EWIF provided by [12] to estimate scope-2 water consumption, as it employs the same methodology for calculating EWIF across a large number of regions. Moreover, a large number of Microsoft’s data centers are located in the U.S., where the average EWIF provided by [12] is 3.14 L/kWh and significantly lower than 4.35 L/kWh reported by the recent U.S. Department of Energy study [5]. The specific location for training GPT-3 is not public. Thus, we consider Microsoft’s different data center locations, while excluding Singapore and Taiwan since the EWIF data for these regions is not available in [12].
3.3.2 Inference
As a representative usage scenario for an LLM, we consider a conversation task, which typically includes a CPU-intensive prompt phase that processes the user’s input (a.k.a., prompt) and a memory-intensive token phase that produces outputs [37]. More specifically, we consider a medium-sized request, each with approximately $\leq$800 words of input and 150 – 300 words of output [37]. The official estimate shows that GPT-3 consumes an order of 0.4 kWh electricity to generate 100 pages of content (e.g., roughly 0.004 kWh per page) [18]. Thus, we consider 0.004 kWh as the per-request server energy consumption for our conversation task. The PUE, WUE, and EWIF are the same as those used for estimating the training water consumption.
Our estimate of inference water consumption for GPT-3 is on the conservative side, and the actual water consumption could be several times higher. Concretely, when service level objectives (SLOs) for LLM response time are considered in real enterprise-grade Nvidia DGX H100 systems for conversation tasks, the inference server energy consumption for a much smaller model (e.g., Llama-3-70B) to process each medium-sized request is already $\sim$0.010 kWh when using a state-of-the-practice LLM inference solution and accounting for the non-GPU server overhead [37]. When considering the Falcon-180B model with a comparable size to GPT-3-175B, the server energy consumption will even reach $\sim$0.016 kWh for processing a medium-sized request [37]. Additionally, our EWIF for the U.S. (i.e., 3.14L/kWh on average) is conservatively chosen and significantly lower than 4.35L/kWh recently reported by [5].
While there is no official information on its resource consumption, the subsequent GPT-4 model is expected to consume substantially more energy and water than GPT-3 for processing the same request [38]. Nonetheless, we emphasize that the on-site WUE of Microsoft’s data centers is already among the lowest in the industry. With continued efforts to reduce AI’s computational demand and improve the overall water efficiency, the water consumption per request may decrease in the future, while the total water consumption is likely to rise further as a result of the growing demand for AI services and the increasing scale of AI applications [5].
4 Our Recommendations
We provide our recommendations to address AI’s water footprint from the scheduling and policy perspectives, making future AI more environmentally sustainable.
(a) Carbon/water efficiency
(b) Hourly carbon/water efficiency
(c) Hourly energy fuel mixes
Figure 2: (a) The U.S. eGRID-level scope-2 water consumption intensity factor vs. carbon emission rate [39, 12]. The dashed line represents a linear regression model, showing that the eGRID-level scope-2 carbon emission and water consumption efficiencies are not aligned. (b) A 5-day snapshot of scope-2 carbon emission rate and water consumption intensity in Virginia, starting from April 4, 2022. The values are calculated based on the fuel mixes, carbon emission rate and water consumption intensity for each fuel type [27, 12, 39]. The scope-2 carbon and water efficiencies only have a weak Pearson correlation coefficient of 0.06 in Virginia. (c) A 5-day snapshot of energy fuel mixes serving Virginia, starting from April 4, 2022 [27].
4.1 More Transparency and Comprehensive Reporting
Despite its growing importance, AI’s water footprint has received relatively less attention. For example, while AI model cards routinely include carbon emissions and serve as an important reporting framework for understanding AI’s environmental impacts, they currently omit information on AI’s water consumption. The lack of transparency may obstruct efforts to drive innovations that enhance water sustainability and support truly sustainable AI. As an initial step to raise awareness among end users about the water resource impacts of their AI usage, we recommend tracking and reporting AI’s water consumption in AI model cards and/or through cloud dashboards.
Moreover, a comprehensive understanding and reporting of AI’s scope-2 water consumption associated with electricity generation remain limited. Although data centers have increasingly adopted climate-conscious cooling system designs to minimize on-site water consumption [8, 26, 24], these efforts primarily focus on scope-1 water usage while largely overlooking scope-2 impacts. Just as addressing scope-2 carbon emissions is important for mitigating climate change, it is equally crucial to address scope-2 water consumption to reduce AI’s “true water cost”, as noted by the U.S. Department of Energy study [5]. To better reflect the true impacts of data centers on water resources, some technology companies such as Meta have begun to include scope-2 water consumption in their sustainability reports [25]. We recommend the reporting of scope-2 water consumption as a standard practice. This approach makes the off-site water consumption visible to AI model developers as well as end users and can unlock new opportunities for demand-side flexibility, thereby alleviating the overall strain on water resources.
Finally, despite the enormous scope-3 supply-chain water footprint [30], there is limited data available for embodied water usage by chip manufacturing. We recommend further research on scope-3 water consumption to achieve a comprehensive understanding of AI’s overall water footprint and to foster corporate water stewardship.
4.2 “When” and “Where” Matter
Judiciously deciding “when” and “where” to train a large AI model can significantly affect the water footprint. The water efficiency exhibits a spatial-temporal diversity — on-site water efficiency changes due to variations of outside weather conditions, and off-site water efficiency changes due to variations of the grid’s energy fuel mixes to meet time-varying demands (Figure 2). Therefore, we can dynamically schedule AI training and inference in a water-wise manner to cut the water footprint. For example, we may schedule AI training at midnight and/or in a data center with better water efficiency. Likewise, if informed of the real-time water efficiency, some water-conscious users may prefer to use AI inference during water-efficient hours and/or in water-efficient data centers, which can reduce AI’s water footprint by enabling demand-side flexibility.
4.3 “Follow the Sun” or “Unfollow the Sun”
To cut the carbon footprint, it is preferable to “follow the sun” when solar energy is more abundant. Nonetheless, to cut the water footprint, it may be more appealing to “unfollow the sun” to avoid high-temperature hours of a day when WUE is high. This conflict can also be shown in Figure 2(a) and Figure 2(b), where we see misalignment between the scope-2 water consumption intensity factor and carbon emission rate: minimizing one footprint might increase the other footprint. This observation further corroborates the previous finding that the environmental impacts of carbon and water footprints are not substitutable [5, 13]. Therefore, to judiciously achieve a balance between “follow the sun” for carbon efficiency and “unfollow the sun” for water efficiency, we need to reconcile the potential water-carbon conflicts by using holistic approaches that are both carbon-efficient and water-wise.
5 Conclusion
In this paper, we uncover AI’s water usage as a critical concern for socially responsible and environmentally sustainable AI. We present a principled methodology to estimate AI’s water footprint. Then, using GPT-3 as an example, we show that a large AI model can consume millions of liters of water for training. We also discuss that the scope-1 and scope-2 water efficiencies vary spatially and temporally — judiciously deciding “when” and “where” to run a large AI model can significantly cut the water footprint. In addition, we recommend increased transparency and comprehensive reporting of AI’s water footprint, and highlight the necessity of holistically addressing the water footprint along with the carbon footprint to build truly sustainable AI.
AI’s water footprint can no longer stay under the radar and must be addressed as a priority as part of the collective efforts to combat global water challenges.
Appendix: Operational Water for Global AI in 2027
A recent study suggests that the global AI could consume 85 – 134 TWh of electricity in 2027 based on the GPU shipment [11], whereas a more aggressive projection by the U.S. Department of Energy predicts that AI servers’ electricity consumption in the U.S. alone will surpass 150 – 300 TWh in 2028 [5]. Based on the former and more conservative projection, we estimate the potential water usage for global AI in 2027, while noting that our global estimates will be exceeded by the water usage attributed to AI in the U.S. alone in 2028 if the U.S. Department of Energy’s projection comes to fruition.
Scope-1 water usage. The scope-1 water efficiency depends on a variety of factors, including the cooling system designs, climate conditions, and operational settings. To set the global scope-1 water efficiency, we utilize the annualized water efficiencies reported by two leading data center operators, Google and Equinix, in their latest sustainability reports [8, 40]. Specifically, for on-site scope-1 water withdrawal, we assume 1.2 L/kWh, which results in a total scope-1 water withdrawal of 0.11 – 0.16 billion cubic meters. Similarly, assuming 1.0 L/kWh for the global scope-1 water consumption efficiency, we obtain a total on-site scope-1 water consumption of 0.09 – 0.14 billion cubic meters. Note that Google and Equinix both operate data centers globally, but represent two distinct categories of data centers: hyperscale data centers (Google) and multi-tenant colocation data centers (Equinix). According to the U.S. Department of Energy [5], these two types of data centers collectively account for the vast majority of data center energy consumption in the U.S., with colocation data centers consuming slightly more energy than hyperscalers.
Scope-2 water usage. As noted by the U.S. Department of Energy [5], scope-2 water usage is part of the true water cost of data centers. The U.S. average electricity water withdrawal and consumption intensity factors are both lower than the global averages [12]. Thus, in our estimate, we use the U.S. average electricity water withdrawal intensity factor 43.83 L/kWh [41], and electricity water consumption intensity factor 3.14 L/kWh [12], respectively. Note that, since [12] includes hydropower in the calculation, it has a higher electricity water withdrawal factor than the U.S. Energy Information Administration’s calculation (i.e., 386.07 L/kWh vs. 43.83 L/kWh for the U.S.). Moreover, our value of 3.14 L/kWh for the U.S. average water consumption factor is lower than 4.35 L/kWh reported by the U.S. Department of Energy [5], as well as lower than Meta’s global electricity water consumption intensity factor of 3.70 L/kWh in 2024 (i.e., 55,475 megaliters divided by 14,975,435 MWh) [25]. Therefore, our choices of 43.83 L/kWh and 3.14 L/kWh for electricity water withdrawal and consumption intensity factors are both on the conservative side, which can partly absorb potential over-estimates of global AI’s energy demand in 2027 provided by [11].
To account for the data center non-IT energy overheads, we conservatively assume a power usage effectiveness (PUE) of 1.1, which is a fairly low value even for state-of-the-art data center facilities [8]. Thus, AI’s total electricity consumption becomes 93.5 – 147.4 TWh. Thus, after multiplying 43.83 L/kWh and 3.14 L/kWh by 93.5 – 147.4 TWh, we obtain the total scope-2 water withdrawal of 4.10 – 6.46 billion cubic meters and water consumption of 0.29 – 0.46 billion cubic meters, respectively.
Total water usage. By adding up scope-1 and scope-2 water usage together, the total water withdrawal and water consumption of global AI may reach 4.2 – 6.6 billion cubic meters and 0.38 – 0.60 billion cubic meters, respectively. According to the U.S. Central Intelligence Agency [42], the estimated U.S. annual water withdrawals in Denmark and the United Kingdom in 2020 (the latest year available as of January, 2025) were 0.98 billion cubic meters and 8.42 billion cubic meters, respectively. Thus, assuming that the 2027 water withdrawals in these two countries remain similar to their 2020 levels, the total water withdrawal attributed to global AI in 2027 is projected to surpass the equivalent of the total annual water withdrawal of 4 – 6 Denmark or approximately half of the United Kingdom. The U.S. Central Intelligence Agency [42] does not provide the country-wide annual water consumption information, and hence we do not contextualize the total water consumption of global AI in 2027.
The estimates of global AI’s water usage in 2027 are naturally subject to uncertainties, e.g., the future water efficiency may differ from the current value we use. Nonetheless, we emphasize that our estimates are on the conservative side. For example, considering a higher estimate, the U.S. Department of Energy’s scope-1 water consumption attributed to AI in the U.S. alone could exceed 0.2 billion cubic meters in 2028 [5]. Moreover, based on the reported scope-2 water consumption efficiency, the combined scope-1 and scope-2 water consumption attributed to AI in the U.S. alone is projected to reach up to about 2 billion cubic meters in 2028 [5], which is significantly higher than our estimate of global AI’s total water consumption in 2027.